I made a voltage divider using a 20.47kohm and 1.07kohm (both values measured with a DMM). The whole thing is packaged inside a Pomona box I found lying around on the Electronics bench. I have hooked it up to the ALSY_I channel and will leave it so overnight. The INMON of this channel isn't DQed, but for this test, the 16Hz EPICS data will suffice. I've locked the EX laser to the arm, enabled slow temperature servo to allow overnight lock (hopefully) and disabled LSC mode (as locking the arm to the MC tends to break the green lock)
To convert the INMON counts to RF power, I will use (based on my earlier calibration of this monitor channel, see DCC document for the demod chassis).

1AM update: Attachment #1 shows that the RF amplitude has been relatively stable (less than 10% of nominal value variation) over the course of the last hour or so. Even though there is some low frequency drift over timescales of ~20mins, no evidence of the wild ~20dB amplitude changes I saw last week. The signs are encouraging...
overnight update: See Attachment #2 - looking at the past 11 hours of second trend data during which the arm stayed locked, there actually seems to have been more significant variation in the beatnote amplitude. Swings of up to 6dBm are seen on a ~20min timescale, while there is also some longer term drift over 12 hours by a couple of dBm. There is probably a systematic error in the Y-axis, as I measured the RF power at the input of the power splitter at the LSC rack to be ~3dBm, so I expect something closer to 0dBm to be the LO input power which is what I am monitoring. So further debugging is required - I think I'll start by aligning the X fiber coupled beam to one of the fiber's special axes. |